OpenStudy (anonymous):

Lincoln and Gemma are looking at the equation the square root of the quantity of 2 times x minus 3 equals square root of x . Lincoln says that the solution is extraneous. Gemma says the solution is non-extraneous. Is Lincoln correct? Is Gemma correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (anonymous):

\[\sqrt{2x-3}=\sqrt{x}\]

OpenStudy (anonymous):

please help me solve this and find out if it extraneous or not :/